OpenStudy (anonymous):

Given the function K described by K(x)=x+1,find each of the following k(0)=?

OpenStudy (anonymous):

plug in zero everywhere you see x

OpenStudy (anonymous):

k(0)=1

OpenStudy (anonymous):

k(0)=0+1

OpenStudy (anonymous):

k(-3)=?

OpenStudy (anonymous):

everywhere you see x plug in -3

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

k(-18)=?

OpenStudy (anonymous):

plug in -18 everywhere you see an x

OpenStudy (anonymous):

hey do just as lagrange said put -18 instead of x f(-18)=? u tell me

OpenStudy (anonymous):

do you see a pattern??

OpenStudy (anonymous):

no I dont

OpenStudy (anonymous):

f(-18)= -18 +1 now tell the answer

OpenStudy (anonymous):

-17?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

k(y+4)=?

OpenStudy (anonymous):

put y+4 instead of x in the eq f(x) = x +1 try and tell the answer :)

OpenStudy (anonymous):

f(y+4)=y+4+1

OpenStudy (anonymous):

5y?

OpenStudy (anonymous):

you have to show me I wont answer u nemore

OpenStudy (anonymous):

then dont

OpenStudy (anonymous):

Im not begging you

OpenStudy (anonymous):

ur not learning , i want u to learn . this is important tc

OpenStudy (anonymous):

put the value instead of x if question say put d in f(x)=x+1 then put d intd of x u will get f(d)=d+1

OpenStudy (anonymous):

I do not understand. But you do not have to help me if you do not want to. I am going into a profession that this will never serve a purpose